A steel railroad rail expands and contracts according to the formula Δl = 𝛼lΔT,

where Δl is the change in length, 𝛼 is a constant called the coefficient of linear expansion, l is the original length, and ΔT is the change in temperature. If a 50.0 ft steel rail is installed at 0°F, how many inches will it expand when the temperature rises to 190°F? (𝛼 = 6.5 10−6/°F.


1
Expert's answer
2022-02-21T14:58:08-0500

We know that

"∆l=l\alpha(∆T)"

"∆l=l

\alpha(T_2-T_1)"

Put value


"∆l=50\times6.5\times10^{-6}(190-0)=0.062ft"